OpenStudy (anonymous):

f(x)=x^2-1; g(t)=sqrt t+1. compute [f(x)]^2

OpenStudy (anonymous):

\[x^4-2 x^2+1 \]

OpenStudy (anonymous):

Le\[f(x)=x^2-1.\]Then\[[f(x)]^2=(x^2-1)^2=x^4-2x^2+1.\]